# Long-Horizon Model-Based Offline Reinforcement Learning Without Conservatism

### v1

@@ G.1 World Model Ensemble and Planning
Planning. At the start of planning, following subsection 4.2, we sample a batch of initial histories $h_{t}\sim\mathcal{D}$, drawn uniformly across time steps. The batch size is fixed at $100$, regardless of the ensemble size $N$. Since the values of $N$ used in our experiments $\{5,20,100\}$ are divisors of $100$, we distribute the histories evenly across ensemble members. The recurrent policy $\pi_{\nu}$ initializes its hidden state $z_{t}$ with $h_{t}$ and then interacts with each world model to generate corresponding rollout until reaching the stopping criterion described in Appendix E. The uncertainty threshold $\mathcal{U}(\zeta)$ for truncation is fixed at $\zeta=1.0$ in our main experiments and varied to $\{0.9,0.99,0.999\}$ in the sensitivity analysis. The entire planning process is parallelized on a GPU and thus introduces only negligible time costs.
@@ G.2 Recurrent Off-Policy RL
Replay buffer size
Full size, i.e.,
(60M in AntMaze, 1M in bandit, 40M otherwise)
@@
Entropy coef. $\alpha$
Auto-tuned with target $\mathrm{dim}(\mathcal{A})$
(AntMaze uses $10\,\mathrm{dim}(\mathcal{A})$ while Hopper uses fixed $\alpha=0.2$)
@@ H.1 Full Benchmarking Results
H.2 Full Results on Compounding Errors
Uncertainty threshold can safeguard against compounding error with LayerNorm. For world ensembles with LayerNorm, we find that the uncertainty threshold $\zeta=1.0$ (used in our main experiments) reliably separates severe error regions, since $\zeta=1.0$ approximates the boundary of in-distribution data. While the Spearman’s rank coefficient, often used to benchmark uncertainty estimation accuracy (Lu et al., 2021), is less than $0.6$ in 5 out of the 18 setups, we argue that this metric is less critical for Bayesian RL. Unlike approaches that penalize with uncertainties at every step, we only use uncertainty as a binary cutoff for truncation. As a result, our method is inherently more robust to imperfect uncertainty ranking, requiring only that severe errors lie beyond the $\zeta=1.0$ boundary.
H.3 Full Results on Adaptive Long-Horizon Planning
@@
A key observation is that using a quantile threshold of $\zeta=0.9$ corresponds roughly to a horizon cap of 111010, which mirrors the short fixed horizons commonly adopted in prior work. This shows that such prior choices are not suitable for guiding the design of Bayesian RL, where adaptive long horizons are essential.

### v2

@@ G.2 Details on Planning
At the start of planning, following subsection 4.2, we sample a batch of initial histories $h_{t}\sim\mathcal{D}$, drawn uniformly across time steps. The batch size is fixed at $100$, regardless of the ensemble size $N$. Since the values of $N$ used in our experiments $\{5,20,100\}$ are divisors of $100$, we distribute the histories evenly across ensemble members. The recurrent policy $\pi_{\nu}$ initializes its hidden state $z_{t}$ with $h_{t}$ and then interacts with each world model to generate corresponding rollout until reaching the stopping criterion described in Appendix E. The uncertainty threshold $\mathcal{U}(\zeta)$ for truncation is fixed at $\zeta=1.0$ in our main experiments and varied to $\{0.9,0.99,0.999\}$ in the sensitivity analysis. The entire planning process is parallelized on a GPU and thus introduces only negligible time costs.
@@ G.3 Details on Recurrent Off-Policy RL
Replay buffer size Full size, i.e., (60M in AntMaze, 1M in bandit, 40M otherwise)
@@
Entropy coef. $\alpha$ Auto-tuned with target $\mathrm{dim}(\mathcal{A})$ (AntMaze uses $10\,\mathrm{dim}(\mathcal{A})$ while Hopper uses fixed $\alpha=0.2$)
@@ H.1 Full Benchmarking Results
H.2 Full Results on Value Overestimation and Horizon Scales
@@
A key observation is that using a quantile threshold of $\zeta=0.9$ corresponds roughly to a horizon cap of 111010, which mirrors the short fixed horizons commonly adopted in prior work. This shows that such prior choices are not suitable for guiding the design of Bayesian RL, where long horizons are essential.
@@
H.3 Full Results on Compounding Errors
Uncertainty threshold can safeguard with LayerNorm against compounding error. For world ensembles with LayerNorm, we find that the uncertainty threshold $\zeta=1.0$ (used in our main experiments) reliably separates severe error regions, since $\zeta=1.0$ approximates the boundary of in-distribution data. While the Spearman’s rank coefficient, often used to benchmark uncertainty estimation accuracy (Lu et al., 2021), is less than $0.6$ in 5 out of the 18 setups, we argue that this metric is less critical for Bayesian RL. Unlike approaches that penalize with uncertainties at every step, we only use uncertainty as a binary cutoff for truncation. As a result, our method is inherently more robust to imperfect uncertainty ranking, requiring only that severe errors lie beyond the $\zeta=1.0$ boundary.